OpenStudy (anonymous):

Suppose A = |a b c| it's a matrix |d e f| |g h i| and det(A) = 15. Using only a simple calculator, find det(M) where M = | a b c | |4d+a 4e+b 4f+c| | g h i |

OpenStudy (anonymous):

who ever want to help me, you can give me steps and I will do it Thanks for everyone!

OpenStudy (shubhamsrg):

det(M)= 1/4 det(A) !

OpenStudy (anonymous):

ok why?

OpenStudy (shubhamsrg):

well the value of determinant dosent change when you add rows simply.. and when you multiply a determinant with some constant,,any row or column gets multiplied with that no..

OpenStudy (shubhamsrg):

i mean on transforming any row,,value of det( ) dosent change. hope i was clear enough.

OpenStudy (anonymous):

yea So the answer is 15/4?

OpenStudy (shubhamsrg):

ohh am sorry,,i meant det(A) = 1/4 det(M)

OpenStudy (shubhamsrg):

ans in 60 then.

OpenStudy (anonymous):

15*4=60?

OpenStudy (shubhamsrg):

the very same..

OpenStudy (anonymous):

yea I didn't see ur answer! lol Thank you so much Can you help me with one more?

OpenStudy (shubhamsrg):

i dont know.. :P

OpenStudy (anonymous):

you don't know what?

OpenStudy (anonymous):

I'm just asking you if you can??

OpenStudy (shubhamsrg):

i dont know if i'd be able to help.. i'll have to see the ques first,,ofcorse..

OpenStudy (anonymous):

ok I will post it as you know it will take time to write a matrix!